56406 is the product of what two consecutive numbers?
Answer by mananth(16946)

let the numbers be x & x+1
..
x(x+1)=56406
x^2+x-56406=0
x^2+238x-237x-56406=0
x(x+238)-237(x+406)=0
(x+238)(x-237)=0
x=-238 OR 237
the numbers are -238, -237
237 & 238